Incomplete manufacturing cost data for Black Company for the year are presented as follows for four different situations:

Direct Materials Used Labour Used Manufacturing Overhead Total manufacturing costs Work in Progress (1/1) Work in Progress (12/31) Cost of Goods Manufactured
1. $127,000 $160,000 $89,000 (a) $33,000 (b) $360,000
2. (C) $180,000 $123,000 $401,000 (d) $40,000 $470,000
3. $80,000 $100,000 (e) $274,000 $60,000 $80,000 (f)
4. $67,000 (g) $75,000 $238,000 $45,000 (h) $270,000

(a) Indicate the missing amount for each letter below. Do not use $ signs or commas when entering your answer:
(a) $

(b) $

(c) $

(d) $

(e) $

(f) $

(g) $

(h) $
Answer #1

Solution

Direct Materials Used Labor Used Manufacturing Overhead Total manufacturing costs Work in Progress (1/1) Work in Progress (12/31) Cost of Goods Manufactured
1 $1,27,000.00 $1,60,000.00 $   89,000.00 $3,76,000.00 $   33,000.00 $   49,000.00 $3,60,000.00
2 $   98,000.00 $1,80,000.00 $1,23,000.00 $4,01,000.00 $1,09,000.00 $   40,000.00 $4,70,000.00
3 $   80,000.00 $1,00,000.00 $   94,000.00 $2,74,000.00 $   60,000.00 $   80,000.00 $2,54,000.00
4 $   67,000.00 $   96,000.00 $   75,000.00 $2,38,000.00 $   45,000.00 $   13,000.00 $2,70,000.00

Working

Direct Materials Used Labor Used Manufacturing Overhead Total manufacturing costs Work in Progress (1/1) Work in Progress (12/31) Cost of Goods Manufactured
1 127000 160000 89000 =127000+160000+89000 33000 =+(376000+33000)-360000 360000
2 =401000-180000-123000 180000 123000 401000 =+(470000+40000)-401000 40000 470000
3 80000 100000 =274000-80000-100000 274000 60000 80000 =274000+60000-80000
4 67000 =238000-67000-75000 75000 238000 45000 =+(238000+45000)-270000 270000